a particle P moves in a straight line . the velocity vms-1 at time t s is given by
v=5t(t-2) for 0<=t<=4
v=k for 4<=t<=14
v=68-2t for 14<=t<=20
where k is constant
1) find k.
2)sketch the velocity time graph for 0<=t<=20
3)find the set of values of t for which the acceleration of P is positive
4)find the total distance travelled by P in the interval 0<=t<=20
1
Expert's answer
2020-04-21T19:06:37-0400
  1. In order to get continuous velocity function it is required "5t(t-2) = k" at "t = 4". Thus "k = 40."

The velocity time graph for 0<=t<=20:

  1. The acceleration of P is positive when velocity increases. According to the graph, it is true for "t\in [1,4]" s.
  2. The total distance is: "d = d_1 + d_2 + d_3" , where:

"d_1 = \int_0^4(5t^2 - 10t) dt\approx26.7\\"

"d_2 = \int_4^{14}40 dt = 400\\"

"d_3 = \int_{14}^{20}(68-2t) dt = 204\\"

Thus, "d = 26.7+400+204=630.7" m.


Answer. 1) k = 40; 2) see graph; 3) t in [1,4]s; 4) d = 630.7m.